$SPAD/src/input f02wef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
m := 5
n := 3
a  : Matrix SF:=
 [[ 2.0, 2.5, 2.5],_
  [ 2.0, 2.5, 2.5],_
  [ 1.6,-0.4, 2.8],_
  [ 2.0,-0.5, 0.5],_
  [ 1.2,-0.3,-2.9] ]
lda := m
ncolb := 1
b : Matrix SF:=
 [[ 1.1, 0.9, 0.6, 0.0, -0.8 ]]
ldb := 5
wantq := true
wantp := true
ldq := 1
ldpt := n
ifail := -1
result:=f02wef(m,n,lda,ncolb,ldb,wantq,ldq,wantp,ldpt,a,b,ifail)

———-

References^{[1] nothing}

4